# Exploring the Potential of Encoder-free Architectures in 3D LMMs

### v1

@@ Abstract
Encoder-free architectures have been preliminarily explored in the 2D visual domain, yet it remains an open question whether they can be effectively applied to 3D understanding scenarios. In this paper, we present the first comprehensive investigation into the potential of encoder-free architectures to overcome the challenges of encoder-based 3D Large Multimodal Models (LMMs). These challenges include the failure to adapt to varying point cloud resolutions and the point features from the encoder not meeting the semantic needs of Large Language Models (LLMs). We identify key aspects for 3D LMMs to remove the encoder and enable the LLM to assume the role of the 3D encoder: 1) We propose the LLM-embedded Semantic Encoding strategy in the pre-training stage, exploring the effects of various point cloud self-supervised losses. And we present the Hybrid Semantic Loss to extract high-level semantics. 2) We introduce the Hierarchical Geometry Aggregation strategy in the instruction tuning stage. This incorporates inductive bias into the LLM early layers to focus on the local details of the point clouds. To the end, we present the first Encoder-free 3D LMM, Enel, whose 7B model rivals the current state-of-the-art ShapeLLM-13B, achieving 55.0%, 50.92%, and 42.7% on the classification, captioning, and VQA tasks, respectively. Our results demonstrate the encoder-free architecture to be highly promising in the field of 3D LMMs. Code is released at https://github.com/Ivan-Tang-3D/ENEL.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023; Bai et al., 2023; Jiang et al., 2023; Cai et al., 2024) have gained unprecedented attention for their proficiency in understanding and generating complex language scenarios. Building upon these advances, many recent efforts have been made to develop Large Multimodal Models (LMMs), empowering LLMs with the capability to interpret multimodal information, such as 2D images (Liu et al., 2024; Li et al., 2024a; Zhang et al., 2024a, c, b; Li et al., 2024b) and 3D point clouds (Guo et al., 2023b; Xu et al., 2025; Guo et al., 2023a; Guo* et al., 2024; Zhang et al., 2022b; Jia et al., 2024).
@@
Specifically for 3D LMMs, the encoder-based architecture has the following potential drawbacks: (1) Point Cloud Resolution Limitation. 3D encoders are often pre-trained on point cloud data at a fixed resolution, such as 1,024 points in PointLLM (Xu et al., 2025). However, during inference, the resolution of point clouds may vary (e.g., 8,192 or 512 points). This difference between training and inference resolutions can result in the loss of spatial information when extracting 3D embeddings, leading to difficulties for LLMs to comprehend, as showcased in Figure 1 (a). (2) Embedding Semantic Discrepancy. 3D encoders are typically pre-trained using self-supervised methods like MAE (Pang et al., 2022; Tang et al., 2024a, b) and contrastive learning (Xie et al., 2020; Qi et al., 2023), but these training objectives may not align with the specific semantic needs of LLMs. In other words, they may not capture the most relevant semantics for LLMs to understand 3D objects, as visualized in Figure 1 (b). Even when a projection layer is used to connect 3D encoders with LLMs, simple MLPs are often insufficient for a complete semantic transformation. Given these issues, we ask: Is it possible to explore an encoder-free architecture for 3D LMMs, eliminating the 3D encoder and instead integrating its functionality directly within the LLM itself?
@@
(1) How can we compensate for the high-level 3D semantics originally extracted by the 3D encoder? In 3D LMMs, the raw point cloud input is first passed through a token embedding module for low-level tokenization, before being processed by the main 3D encoder, usually a Transformer (Vaswani, 2017), to generate high-level embeddings. Skipping the encoder entirely poses a challenge in capturing the complex spatial structures of 3D point clouds. To address this, we propose a strategy called LLM-embedded Semantic Encoding in the pre-training stage. First, we adopt a simple yet effective token embedding module that captures as much informative semantic content as possible. These 3D tokens are then directly fed into the LLM. Next, we aim to shift the responsibility of capturing high-level 3D semantics to the LLM itself. To facilitate this, we make the early layers of the LLM learnable, allowing them to specialize in 3D encoding. To guide this process, we explore various 3D self-supervised loss functions, such as reconstruction loss, masked modeling loss, and distillation loss, and ultimately propose the Hybrid Semantic Loss as the most effective choice.
@@
(2) How can we integrate inductive bias into LLMs for better perception of 3D geometric structures? Traditional 3D encoders typically embed explicit inductive bias into their architectures to progressively capture multi-level 3D geometries. For instance, models like Point-M2AE (Zhang et al., 2022a) use a local-to-global hierarchy, which is a concept also common in convolutional layers for 2D image processing (He et al., 2016). In contrast, LLMs employ standard Transformer architectures, where each layer processes the same number of tokens, representing the same semantic level across the network. In the absence of the encoder, we introduce the approach of Hierarchical Geometry Aggregation during the second fine-tuning stage. In the early layers of the LLM, we aggregate 3D tokens based on their geometric distribution using Farthest Point Sampling (FPS) and $k$-Nearest Neighbor ($k$-NN) sampling. This approach enables the LLM to gradually integrate detailed 3D semantics and develop a more holistic understanding of the 3D object. In the later layers, we reverse this aggregation, propagating the tokens back to their original distribution to maintain the fine-grained representation necessary for effective semantic communication.
@@
Through a series of experimental investigations, we have uncovered the strong potential of applying encoder-free architecture to the 3D LMM domain. Building on our insights, we introduce Enel, an ENcoder-freE 3D LMM evolved from Vicuna-7B (Chiang et al., 2023) using the same training dataset from PointLLM. Notably, without any 3D encoders, Enel achieves comparable performance to the current state-of-the-art ShapeLLM-13B (Qi et al., 2024), attaining scores of 55.0% and 50.92% on the classification and captioning tasks, respectively. We hope Enel may provide the community with a scalable and effective path for adapting the encoder-free architecture to 3D scenarios.
@@
• We further introduce Enel, a concise and well-performed encoder-free 3D LMM, which achieves 50.92%, 55.0% and 42.70% on 3D captioning, classification, and 3D VQA tasks, respectively, on par with existing encoder-based models.
@@ 2.1 Overall Architecture
We present the first attempt to extend the encoder-free architecture to 3D LMMs, such as PointLLM (Xu et al., 2023) and ShapeLLM (Qi et al., 2024), in order to efficiently handle the complex tasks like embodied agent (Guo et al., 2023a) and vision-language navigation. We select PointLLM as the baseline model for the exploration and evaluate the performance of different strategies on the Objaverse dataset (Deitke et al., 2023), using GPT-4 scores combined with traditional metrics as our evaluation metrics. The benchmark is highly challenging, as it requires the model to achieve high-quality alignment between 3D semantics and textual space while also capturing the intricate geometric structures of 3D objects.
In the overall architecture, the encoder-free 3D LMM directly utilizes a token embedding module to convert point cloud data into discrete point tokens, which are then concatenated with text tokens to serve as input to the LLM. As shown in Figure 2, to assume the role of the encoder, the LLM is guided to extract high-level semantic features of the point clouds and acquire multi-level knowledge from both global and local perspectives. In the subsequent sections, we primarily explore two strategies within the encoder-free architecture: LLM-embedded Semantic Encoding (Section 2.2) and Hierarchical Geometry Aggregation (Section 2.3).
Table 1: Token Embedding. We evaluate the performance on the Objaverse benchmark and adopt PointLLM-7B as the baseline model. ’Cls’ and ’Cap’ represent classification and captioning tasks, respectively. S-BERT refers to the Sentence-BERT. T.E. stands for our designed token embedding module.
Method Cls Cap
GPT-4 GPT-4 S-BERT
PointLLM-7B 53.00 44.85 47.47
- Encoder 35.50 33.37 41.19
+ 2-layer T.E. 42.50 41.35 44.25
+ 3-layer T.E. 47.31 43.86 45.89
+ 4-layer T.E. 45.00 42.99 44.51
@@
Further 3D Encoding. We discover that the absence of the encoder results in a lack of context modeling in point cloud feature processing. Therefore, we attempt to have the early layers of the LLM take on the encoder’s role in capturing global interactions of features, further encoding the point cloud features. In the pre-training stage, we set the first K layers of the frozen LLM to be learnable, utilizing the self-attention mechanism to capture global geometric structures. Meanwhile, we experiment with both the original learning rate and a smaller learning rate. As shown in Table 2, the smaller learning rate generally leads to better results. This is because a smaller learning rate can make the optimization process of the early layers more stable. Based on the designed token embedding module, setting the first four layers to be learnable yields the best results, as it effectively encodes low-level features into high-level representations with considerable computational efficiency.
Table 2: Further 3D Encoding. We set the LLM early layers to be learnable. LR represents the learning rate during the pre-training stage, with the original learning rate set to 2e-3.
Method LR Cls Cap
GPT-4 GPT-4 S-BERT
PointLLM-7B 2e-3 53.00 44.85 47.47
+ 2 learnable layers 2e-3 41.06 42.23 45.92
4e-4 45.5 44.72 47.35
+ 4 learnable layers 2e-3 44.85 41.53 46.77
4e-4 49.11 45.39 47.71
+ 8 learnable layers 2e-3 43.76 39.71 42.38
4e-4 48.00 44.49 47.21
2.2 LLM-embedded Semantic Encoding
Masked Modeling Loss. In the pre-training stage, we apply the Masked Modeling Loss to the point tokens processed by the LLM, as shown in Figure 3 (a). Through the Farthest Point Sampling (FPS) and k-Nearest Neighbors (k-NN) algorithms in the token embedding, the point clouds $\{P_{i}\}_{i=1}^{N}$ are divided into point patches $\{G_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times k\times 3}$ and the corresponding point tokens $\{F_{i}\}_{i=1}^{M}$. We randomly mask the point tokens with a masking ratio $r$, and replace them with learnable tokens. The masked feature tokens can be denoted as $\{F_{\text{gt}_{i}}\}_{i=1}^{M*r}$, which serve as the ground truth for the loss computation. After the learnable tokens are concatenated with visible tokens and processed by the LLM, a linear layer is applied to extract the point tokens $\{F_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times D_{1}}$, and the Mean Squared Error (MSE) is computed between the predicted $F_{\text{pre}}$ and the ground truth $F_{\text{gt}}$. The optimization can be written as
The specific process of applying Masked Modeling to point patches $G$ is detailed in Appendix C.1.
where $a=G_{\text{pre}}$ and $b=G$. The detailed procedure for reconstructing feature tokens F can be found in Appendix C.1.
Table 3: LLM-embedded Semantic Encoding. In the pre-training stage, we explore the effects of various self-supervised learning losses targeting point tokens. $\Psi$ represents a mask ratio of 60%, while $\Phi$ represents a mask ratio of 30%. The subscript patch and feat represent the loss target. For Hybrid Semantic Loss, the subscript patch and feat represent the masked modeling target, while the reconstruction target is the corresponding feat and patch.
Method Cls Cap
@@ 2.2 LLM-embedded Semantic Encoding
[…]
@@
[…]
+ Position Embedding 53.00 48.85 48.00
Contrastive Loss. We conduct contrastive learning (Khosla et al., 2020) at the point cloud level, where we contrast two transformed versions of the point cloud in the Figure 3 (c). Given a sampled point cloud $\{P_{i}\}_{i=1}^{N}$, we apply two random geometric transformations $T_{1}$ and $T_{2}$, including rotation and translation, to obtain $P_{T1}$ and $P_{T2}$. The two augmented point clouds are separately paired with the original text query and processed through the LLM to obtain their respective feature tokens $F_{T1}\in\mathbb{R}^{M\times D_{1}}$ and $F_{T2}\in\mathbb{R}^{M\times D_{1}}$. Within the mini-batch, the two feature tokens derived from the same point cloud serve as positive pairs, while they are considered negative pairs with other point clouds. Based on the NCESoftmaxLoss, we aim to maximize the similarity of positive pairs and minimize the similarity of negative pairs, forcing the LLM to learn the geometric equivariance of point clouds. The specific formula $\mathcal{L}_{\text{contrast}}$ is as follows:

### v2

@@ Abstract
Encoder-free architectures have been preliminarily explored in the 2D visual domain, yet it remains an open question whether they can be effectively applied to 3D understanding scenarios. In this paper, we present the first comprehensive investigation into the potential of encoder-free architectures to alleviate the challenges of encoder-based 3D Large Multimodal Models (LMMs). These challenges include the failure to adapt to varying point cloud resolutions and the point features from the encoder not meeting the semantic needs of Large Language Models (LLMs). We identify key aspects for 3D LMMs to remove the encoder and enable the LLM to assume the role of the 3D encoder: 1) We propose the LLM-embedded Semantic Encoding strategy in the pre-training stage, exploring the effects of various point cloud self-supervised losses. And we present the Hybrid Semantic Loss to extract high-level semantics. 2) We introduce the Hierarchical Geometry Aggregation strategy in the instruction tuning stage. This incorporates inductive bias into the LLM layers to focus on the local details of the point clouds. To the end, we present the first Encoder-free 3D LMM, Enel. Our 7B model rivals the current state-of-the-art model, ShapeLLM-13B, achieving 55.10%, 50.98%, and 43.10% on the classification, captioning, and VQA tasks, respectively. Our results demonstrate that the encoder-free architecture is highly promising for replacing encoder-based architectures in the field of 3D understanding. Code is released at https://github.com/Ivan-Tang-3D/ENEL.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023; Bai et al., 2023; Jiang et al., 2023; Cai et al., 2024) have gained unprecedented attention for their proficiency in understanding and generating complex language scenarios. Building upon these advances, many recent efforts have been made to develop Large Multimodal Models (LMMs), empowering LLMs with the capability to interpret multimodal information 2D images (Liu et al., 2024; Li et al., 2024) and 3D point clouds (Guo et al., 2023b; Xu et al., 2025), as well as generate multimodal asserts (Xie et al., 2024; Chen et al., 2025; Guo et al., 2025; Jiang et al., 2025).
@@
Specifically for 3D LMMs, the encoder-based architecture has the following potential drawbacks: (1) Point Cloud Resolution Limitation. 3D encoders are often pre-trained on point cloud data at a fixed resolution, such as 8,192 points for Point-BERT (Yu et al., 2022) in PointLLM (Xu et al., 2025). However, during inference, the resolution of point clouds may vary (e.g., 12,000 or 4,000 points). This difference between training and inference resolutions can result in the loss of spatial information when extracting 3D embeddings, leading to difficulties for LLMs to comprehend, as showcased in Figure 1 (a). (2) Embedding Semantic Discrepancy. 3D encoders are typically pre-trained using self-supervised methods like MAE (Pang et al., 2022; Tang et al., 2024a, b) and contrastive learning (Xie et al., 2020; Qi et al., 2023), but these training objectives may not align with the specific semantic needs of LLMs. In other words, they may not capture the most relevant semantics for LLMs to understand 3D objects, as visualized in Figure 1 (b). Even when a projection layer is used to connect 3D encoders with LLMs, simple MLPs are often insufficient for a complete semantic transformation. Given these issues, we ask: Is it possible to explore an encoder-free architecture for 3D LMMs, eliminating the 3D encoder and instead integrating its functionality directly within the LLM itself?
@@
(1) How can we compensate for the high-level 3D semantics originally extracted by the 3D encoder? In 3D LMMs, the raw point cloud input is first passed through a token embedding module for low-level tokenization, before being processed by the main 3D encoder, usually a Transformer (Vaswani, 2017), to generate high-level embeddings. Skipping the encoder entirely poses a challenge in capturing the complex spatial structures of 3D point clouds. To address this, we propose a strategy called LLM-embedded Semantic Encoding in the pre-training stage. First, we adopt a simple yet effective token embedding module that captures as much informative semantic content as possible. These 3D tokens are then directly fed into the LLM. Next, we aim to shift the responsibility of capturing high-level 3D semantics to the LLM itself. To facilitate this, we make the early layers of the LLM learnable, allowing them to specialize in multimodal alignment and further 3D encoding. To guide this process, we explore various 3D self-supervised loss functions, such as masked modeling loss and distillation loss, and ultimately propose the Hybrid Semantic Loss as the most effective choice.
@@
(2) How can we integrate inductive bias into LLMs for better perception of 3D geometric structures? Traditional 3D encoders typically embed explicit inductive bias into their architectures to progressively capture multi-level 3D geometries. For instance, models like Point-M2AE (Zhang et al., 2022) use a local-to-global hierarchy, which is a concept also common in convolutional layers for 2D image processing (He et al., 2016). In contrast, LLMs employ standard Transformer architectures, where each layer processes the same number of tokens, representing the same semantic level across the network. In the absence of the encoder, we introduce the approach of Hierarchical Geometry Aggregation during the second fine-tuning stage. In the early layers of the LLM, we aggregate 3D tokens based on their geometric distribution using Farthest Point Sampling (FPS) and $k$-Nearest Neighbor ($k$-NN) sampling. This approach enables the LLM to gradually integrate detailed 3D semantics and develop a more holistic understanding of the 3D object. In the later layers, we reverse this aggregation, propagating the tokens back to their original distribution to maintain the fine-grained representation necessary for effective semantic communication.
@@
Through a series of experimental investigations, we have uncovered the strong potential of applying encoder-free architecture to the 3D LMM domain. Building on our insights, we introduce Enel, an ENcoder-freE 3D LMM evolved from Vicuna-7B (Chiang et al., 2023) using the same training dataset from PointLLM. Notably, without any 3D encoders, Enel-7B achieves comparable performance to the current state-of-the-art ShapeLLM-13B (Qi et al., 2024), attaining scores of 55.10% and 50.98% on the classification and captioning tasks, respectively. We hope Enel may provide the community with a scalable and effective path for adapting the encoder-free architecture to 3D scenarios.
@@
• We further introduce Enel, a concise and well-performed encoder-free 3D LMM, which, at the 7B parameter scale, achieves 50.98%, 55.10% and 43.10% on 3D captioning, classification, and 3D VQA tasks, respectively, on par with existing encoder-based models.
Table 1: Token Embedding. Performance on Objaverse with PointLLM-7B as the baseline. ‘Cls’/‘Cap’: classification/captioning tasks. ‘Avg’: accuracy under prompts “What is this?” and “This is an object of.” ‘S-BERT’: Sentence-BERT. ‘T.E.’: our designed token embedding module.
Method Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 53.00 44.85 47.47 - Encoder 35.50 33.37 41.19 + 2-layer T.E. 42.50 41.35 44.25 + 3-layer T.E. 47.31 43.86 45.89 + 4-layer T.E. 45.00 42.99 44.51
Table 2: Learnable Layers. We set the LLM early layers to be learnable. ‘LR’ represents the learning rate during the pre-training stage, with the original learning rate set to 2e-3.
Method LR Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 2e-3 53.00 44.85 47.47 + 2 learnable layers 2e-3 41.06 42.23 45.92 4e-4 45.50 44.72 47.35 + 4 learnable layers 2e-3 44.85 41.53 46.77 4e-4 49.11 45.39 47.71 + 8 learnable layers 2e-3 43.76 39.71 42.38 4e-4 48.00 44.49 47.21 + 12 learnable layers 2e-3 42.91 40.05 42.40 4e-4 47.63 44.20 47.09
2 Related Work
Encoder-free Vision-Language Models. Traditional vision-language models (VLMs) often rely on vision encoders to extract visual features before processing them with language models, integrating image encoders like CLIP (Radford et al., 2021) and DINO V2 (Oquab et al., 2023). However, recent efforts have explored encoder-free VLMs for their simplicity. Approaches like (ChameleonTeam, 2024; Xie et al., 2024) use VQ tokenizers (Esser et al., 2021) or linear projection layers (Diao et al., 2024a; Chen et al., 2024a) to represent images. Fuyu-8B (Bavishi et al., 2023), a pure decoder-only model, directly processes image patches through linear projections, handling high-resolution images but showing only average performance. The EVE series (Diao et al., 2024b) eliminates the need for a separate vision encoder by bridging vision-language representation within a unified decoder and enhancing visual recognition capabilities through additional supervision.
@@ 3.1 Overall Architecture
We present the first attempt to extend the encoder-free architecture to 3D LMMs, such as PointLLM (Xu et al., 2023) and ShapeLLM (Qi et al., 2024), in order to efficiently handle the complex tasks like embodied agent (Guo et al., 2023a). We select PointLLM as the baseline model for the exploration and evaluate the performance of different strategies on the Objaverse dataset (Deitke et al., 2023), using GPT-4 scores combined with traditional metrics as our evaluation metrics. The benchmark poses a significant challenge, requiring the model to align 3D semantics with text while capturing complex 3D geometries.
In the overall architecture, the encoder-free 3D LMM directly utilizes the token embedding module to convert point clouds into discrete point tokens, which are then concatenated with text tokens to serve as input to the LLM. As shown in Figure 2, to assume the role of the encoder, the LLM is guided to extract high-level semantic features of the point clouds and acquire multi-level knowledge from both global and local perspectives. In the subsequent sections, we primarily explore two strategies within the encoder-free architecture: LLM-embedded Semantic Encoding (Section 3.2) and Hierarchical Geometry Aggregation (Section 3.3).
@@
Table 3: LLM-embedded Semantic Encoding. In pre-training, we explore the effects of different self-supervised learning losses targeting point tokens. $\Psi$ and $\Phi$ denote mask ratios of 60% and 30%, respectively. Subscripts $patch$ and $feat$ indicate loss targets. For Hybrid Semantic Loss, the subscripts $patch$ and $feat$ refer to the masked modeling target, with reconstruction targeting the corresponding $feat$ and $patch$.
Method Cls (Avg) Cap
@@
Masked Modeling LossfeatΨ 50.00 45.80 46.29
Masked Modeling LossfeatΦ 49.50 47.35 47.93
Reconstruction Losspatch 49.50 46.96 47.33
Reconstruction Lossfeat 48.50 45.95 47.18
@@
Hybrid Semantic Losspatch 50.50 46.84 47.59
Hybrid Semantic Lossfeat 53.00 48.85 48.00
Further 3D Encoding & Alignment. We discover that the absence of the encoder results in a lack of context modeling in point cloud feature processing. Therefore, we attempt to have the early layers of the LLM take on the encoder’s role in capturing global interactions of features, further encoding the point cloud features. In the pre-training stage, we set the first K layers of the frozen LLM to be learnable, utilizing the self-attention mechanism to capture global geometric structures. In the first K layers, 3D tokens and text tokens interact and align naturally within a shared semantic space. Meanwhile, we experiment with different learning rates. As shown in Table 1, a smaller learning rate yields better results by stabilizing early layer optimization. Based on the designed token embedding module, setting the first four layers to be learnable yields the best results, as it effectively encodes low-level features into high-level representations with considerable computational efficiency.
3.2 LLM-embedded Semantic Encoding
Masked Modeling Loss. In the pre-training stage, we apply the Masked Modeling Loss to the point tokens processed by the LLM, as shown in Figure 3 (a). Through the Farthest Point Sampling (FPS) and k-Nearest Neighbors (k-NN) algorithms in the token embedding, the point clouds $\{P_{i}\}_{i=1}^{N}$ are divided into point patches $\{G_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times k\times 3}$ and the corresponding point tokens $\{F_{i}\}_{i=1}^{M}$. We randomly mask the point tokens with a masking ratio $r$, and replace them with learnable tokens. The masked feature tokens can be denoted as $\{F_{\text{gt}_{i}}\}_{i=1}^{M*r}$, which serve as the ground truth for the loss computation. After the learnable tokens are concatenated with visible tokens and processed by the LLM, a linear layer is applied to predict the point tokens $\{F_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times D_{1}}$, and the Mean Squared Error (MSE) is computed between $F_{\text{pre}}$ and $F_{\text{gt}}$. The optimization can be written as
The specific process of applying Masked Modeling to point patches $G$ is detailed in Appendix B.1.
where $a=G_{\text{pre}}$, $b=G$. The procedure for reconstructing feature F is detailed in Appendix B.1.
Contrastive Loss. We conduct contrastive learning (Khosla et al., 2020) at the point cloud level, where we contrast two transformed versions of the point cloud in the Figure 3 (c). Given a sampled point cloud $\{P_{i}\}_{i=1}^{N}$, we apply two random geometric transformations $T_{1}$ and $T_{2}$, including rotation and translation, to obtain $P_{T1}$ and $P_{T2}$. The two augmented point clouds are separately paired with the original text query and processed through the LLM to obtain their respective feature tokens $F_{T1}\in\mathbb{R}^{M\times D_{1}}$ and $F_{T2}\in\mathbb{R}^{M\times D_{1}}$. Within the mini-batch, the two feature tokens derived from the same point cloud serve as positive pairs, while they are considered negative pairs with other point clouds. Using NCESoftmaxLoss, we aim to maximize the similarity of positive pairs and minimize the similarity of negative pairs, encouraging the LLM to learn geometric equivariance of point clouds. The $\mathcal{L}_{\text{contrast}}$ is: